Suppose we have a 12.2-L sample containing 0.50 mol oxygen gas (O2) at a pressure of 1 atm and a temperature of 25ºC. If all this O2 were converted to ozone (O3) at the same temperature and pressure, what would be the volume of the ozone

1
Expert's answer
2021-03-05T03:33:29-0500

3O2---> 2O3

3moles of oxygen gives 2moles of O3

0.5moles of oxygen will give 2/3 x 0.5= 0.33mol

P=1atm, V=?, n=0.33mol, R= 0.0821L.atm/K.mol, T= 298K

From ideal gas equation,

PV=nRT

V= nRT/P

V= 0.33 x 0.0821 x 298/1= 8.15L